Plane A travels at 200 mph and plane B travels at 500 mph. If plane A has a 2 hours head start how far will both planes travel before plane B overtakes plane A? The book gives me the answer, I need to see the equation please.
Answer by Alan3354(30993)

Plane A travels at 200 mph and plane B travels at 500 mph. If plane A has a 2 hours head start how far will both planes travel before plane B overtakes plane A? The book gives me the answer, I need to see the equation please.
-----------
A is 400 miles away (200*2)
B gains on A at 300 mph (500 - 200)
B overtakes A in 4/3 hours (400/300)
d = r*t = (4/3)*500
d = 2000/3 miles